Really need questions 6 and 7 answered...
8-37Activity-based costing, batch-level varlance analysis. Rae Steven Publishing Company specializes in printing specialty textbooks for a small but profitable college market. Due to the high setup costs for each batch printed, Rae Steven holds the book requests until demand for a book is approximately 520. At that point Rae Steven will schedule the setup and production of the book. For rush orders, Rae Steven will produce smaller batches for an additional charge of $987 per setup. Budgeted and actual costs for the printing process for 2014 were as follows: Static-Budget Amounts Actual Results Number of books produced 197,600 225,680 Average number of books per setup 520 496 Hours to set up printers 7 hours 7.5 hours Direct variable cost per setup-hour $130 $70 $53,200 68,000 Total fixed setup overhead costs Required
1. What is the static budget number of setups for 20147 2. What is the flexible budget number of setups for 2014? 3. What is the actual number of setups in 2014? 4. Assuming fixed setup overhead costs are allocated using setup-hours, what is the predetermined fixed setup overhead allocation rate? 5. Does Rae Stevens charge of $987 cover the budgeted direct variable cost of an order? The budgeted total cost? 6. For direct variable setup costs, computethe price and efficiency variances. 7. For fixed setup overhead costs, compute the spending and the production-volume variances. 8. What qualitative factors should Rae Steven consider before accepting or rejecting a special order?
Answer #1

1. Static budget number of setups = Budgeted books produced/ Budgeted books per setup= 197,600 ÷ 520 = 380 setups

2. Flexible budget number of setups = Actual books produced / Budgeted books per setup= 225,680 ÷ 520 = 434 setups

3. Actual number of setups = Actual books produced / Actual books per setup= 225,680/496 = 455 setups

4. Static budget number of hours = Static budget number of setups × Budgeted hours per setup= 380 × 7 = 2,660 hours

Fixed overhead rate = Static budget fixed overhead / Static budget number of hours= $53,200/2,660 = $20 per hour

5. Budgeted direct variable cost of a setup = Budgeted variable cost per setup-hour × Budgeted number of setup-hours= $130 × 7 = $910.

Budgeted total cost of a setup = Budgeted direct variable cost + (Fixed overhead rate × Budgeted number of setup-hours)= $910 + ($20 × 7) = $1,050.

So, the charge of $987 covers the budgeted incremental (i.e., variable) cost of a setup but not the budgeted full cost.

6. Direct Variable Variance Analysis for Rae Steven Publishing Company for 2014 :

Actual Variable Cost = 455 * 7.5 * $70 = $238,875

Actual Hours * Budgeted Rate = 455 * 7.5 * $130 = $443,625

Standard Hours * Standard Rate = 434 * 7.0 * $130 = 394,940.

So, here the Price Variance = Actual Variable Cost - (Actual Hours * Budgeted Rate) = $238,875 - $443,625 = $204,750 F

Effeciency Variance = (Actual Hours * Budgeted Rate) - (Standard Hours * Standard Rate) = $443,625 - $394,940 = $48,685U

7. Fixed Setup Overhead Variance Analysis for Rae Steven Publishing Company for 2014

Actual Fixed Overhead = $68,000

Static Budget Fixed Overhead = $53,200

Standard Hours * Budgeted Rate = 434 * 7.0 * $20 = $60,760

Spending Variance = Actual Fixed Overhead - Static Budget Fixed Overhead = $68,000 - $53,200 = $14,800 U

Production Volume Variance = Static Budget Fixed Overhead - (Standard Hours * Budgeted Rate) = $53,200 - $60,760 = $7,560F